OpenStudy (anonymous):

Explain, in complete sentences, how you would factor 8x6 – 56x3 + 98. Provide the factors.

OpenStudy (lgbasallote):

i would take out 2 first...

OpenStudy (anonymous):

could someone show me on the drawing thing.

OpenStudy (anonymous):

8x^6-56x^3+98= 2(4x^6-28x^3+49) let u=x^3 2(4u^2-28u+49) 2(2u-7)^2 put x^3 back in for u 2(2x^3-7)^2